### v7

# Outer approximations of core points for integer programming

Abstract
For several decades the dominant techniques for integer linear programming have been branching and cutting planes. Recently, several authors have developed core point methods for solving symmetric integer linear programs (ILPs). An integer point is called a core point if its orbit polytope is lattice-free. It has been shown that for symmetric ILPs, optimizing over the set of core points gives the same answer as considering the entire space. Existing core point techniques rely on the number of core points (or equivalence classes) being finite, which requires special symmetry groups. In this paper we develop some new methods for solving symmetric ILPs — based on outer approximations of core points — that do not depend on finiteness but are more efficient if the group has large disjoint cycles in its set of generators.
1 Introduction
Formulation symmetries occur in practice when relabellings yield equivalent problem structure; this causes repeated work for branching solvers, and state of the art commercial and research solvers make efforts to break symmetries PR2019 . Let $G\leqslant S_{n}$ be a permutation group acting on $\mathbb{R}^{n}$ by permuting coordinates. For any integer point $z\in\mathbb{Z}^{n}$, the orbit polytope of $z$ is the convex hull of the $G$-orbit of $z$. If the vertices of an orbit polytope are the only integer points in the polytope we call it lattice-free and call $z$ a core point. Instead of seeing symmetry as a problem, core point techniques seek to exploit it to solve integer linear programs (ILPs) faster. In the most direct approach, when the number of core points is finite (which only holds for certain special groups), core points are enumerated and tested individually BHJ2013 . It should be noted that core point techniques are not useful for binary problems since every $\{0,1\}$-point is a core point; HP2019 considers an alternative approach based on lexicographical order.
Computation of symmetries in the MIPLIB 2010 and 2017 instances has been done in miplib and this study shows that many instances are affected by symmetry. Symmetric Integer Linear Programming appears in many problems such as scheduling on identical machines and code construction. For solving these problems, Artificial Intelligence approaches have been investigated in AP1 ; AP2 .
A (not necessarily polyhedral) outer approximation is a set of constraints that is feasible for all of the points in the set one wishes to approximate. A well known example of an outer approximation is an ILP, where the (initial) linear constraints define an outer approximation of the feasible integer points. Outer approximations lead naturally to a hybrid approach where synthesized constraints are added to an existing formulation and then solved with a traditional solver. Outer approximations are implicit in previous results bounding the distance of core points to certain linear subspaces (see e.g. Theorem 3.24 in 2 ). The distance bounds do not themselves seem to be tight enough to provide a practical improvement for solving ILPs. In this paper we develop some new constraints for problems with formulation symmetries. While these constraints are nonlinear and non-convex, initial experiments with nonlinear solvers seem promising.
In Section 2 we give some basic definitions. In Section 3 we consider integer linear programs with cyclic symmetry groups and provide some new constraints to determine outer approximations of their core points. We also provide an algorithm that uses these constraints to solve an ILP. In Section 4 we generalize this algorithm for ILPs where only some of their variables have cyclic symmetry. In Section 5 we generalize the algorithm for direct products of cyclic groups. In Section 6 we classify permutation groups based on their generators and explain how the algorithms of the previous sections can be applied to ILPs having arbitrary permutation groups as symmetry groups. Finally in Section 7 we use our methods to solve some hard symmetric integer linear feasibility problems.
2 Basic Definitions
Symmetries of geometric objects (e.g. polyhedra and integer lattices) in integer linear programming can be viewed as the action of some underlying group.
Definition 1
If $G$ is a group and $X$ is a set, then a (left) group action $\phi$ of $G$ on $X$ is a function
$\phi:G\times X$ $\rightarrow X,$
$(g,x)$ $\rightarrow gx,$
that satisfies the following two axioms.
  1. 1.
    ex = x for all $x\in X$, where $e\in G$ is the identity element.
  2. 2.
    (gh)x = g(hx) for all $g,h\in G$ and all $x\in X$.
In this paper we assume that all groups considered are permutation groups, and that they act on a family $\mathbb{X}^{n}$ of $n$-tuples (e.g. $\mathbb{C}^{n}$, $\mathbb{R}^{n}$, or $\mathbb{Z}^{n}$ in the usual coordinates) as follows.
Remark 1
[…]
$\phi_{g}(x)=(x_{g^{-1}(0)},x_{g^{-1}(1)},\ldots,x_{g^{-1}(n-1)}).$
There are certainly more general notions of symmetry possible, but permuting coordinates is the most widely studied class of symmetries in integer linear programming PR2019 .
The techniques in this paper rely on restricting the search for feasible integral points to certain special integral points called core points. Core points are defined relative to a given group as follows.
Definition 2 (Core Points)
  1. 1.
    Let $P\subset\mathbb{R}^{n}$ be a convex polytope with integral vertices. We call $P$ lattice-free if $P\cap\mathbb{Z}^{n}=\mathrm{vert}(P)$ where $\mathrm{vert}(P)$ is the set of vertices of $P$.
  2. 2.
    Let $G\leqslant GL_{n}(\mathbb{R})$ be a finite group and let $G_{z}$ be the G-orbit of some point $z\in\mathbb{\mathbb{R}}^{n}$. We call the convex hull of this orbit an orbit polytope and denote it by $\mathrm{conv}(G_{z})$.
  3. 3.
    Let $G\leqslant GL_{n}(\mathbb{Z})$ be a finite group of unimodular matrices. A point $z\in\mathbb{Z}^{n}$ is called a core point for $G$ if and only if the orbit polytope $\mathrm{conv}(G_{z})$ is lattice-free.
Let $C_{n}=\langle\sigma\rangle$ denote the cyclic group generated by $\sigma$, a cyclic permutation of coordinates. In other words, $\sigma:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ is given by
$\sigma(c_{0},c_{1},\ldots,c_{n-1})=(c_{n-1},c_{0},\ldots,c_{n-2})$
We represent points in $\mathbb{R}^{n}$ by column vectors but for convenience, we write such vectors here in a transpose way. The map $\sigma$ is easily iterated:
$\sigma^{k}(c_{0},c_{1},\ldots,c_{n-1})=(c_{n-k},c_{n-k+1},\ldots,c_{n-1},c_{0},\ldots,c_{n-k-1})$
Consider a cyclic group $C_{4}=\langle(1,2,3,4)\rangle$. The orbit of $x=(2,0,4,1)$ consists of the four vectors $x_{1}=(2,0,4,1)$, $x_{2}=(1,2,0,4)$, $x_{3}=(4,1,2,0)$, $x_{4}=(0,4,1,2)$. The orbit polytope of $x$ is not lattice-free because
$(1,2,1,3)=\frac{1}{5}x_{1}+\frac{3}{5}x_{2}+0x_{3}+\frac{1}{5}x_{4},$
so that $x$ is not a core point.
Definition 3
Let $G$ act on $\mathbb{C}^{n}$. The subset of $\mathbb{C}^{n}$ preserved by all elements of $G$ is called the fixed space:
$\mathrm{Fix}(G):=\{x\in\mathbb{C}^{n}\;\;|\;\;gx=x,\forall g\in G\}.$
We denote by $\mathrm{Fix}_{\mathbb{Z}}(G)$ (resp. $\mathrm{Fix}_{\mathbb{R}}(G)$), the intersection $\mathrm{Fix}(G)\cap\mathbb{Z}^{n}$ (resp. $\mathrm{Fix}(G)\cap\mathbb{R}^{n}$).
We use $\mathbf{1}^{k}$ to denote the $k$-dimensional vector of all ones (or just $\mathbf{1}$ where the dimension is clear from context). It is easy to see that $\mathbf{1}^{n}$ is contained in $\mathrm{Fix}_{\mathbb{Z}}(G)$ for any group $G$ acting on $\mathbb{R}^{n}$ by permuting coordinates. If $G$ is further transitive, then $\mathrm{Fix}_{\mathbb{R}}(G)=\mathrm{Span}(\mathbf{1})$.
Definition 4
We define the $k$-th layer to be the set
$\mathbb{Z}_{(k)}^{n}:=\{z\in\mathbb{Z}^{n}\;\;|\;\;\langle z,\mathbf{1}\rangle=k\}.$
Note that the set $\mathbb{Z}_{(k)}^{n}$ is $G$-invariant because $G$ acts by permuting coordinates. Consider the cyclic group $C_{4}=\langle(1,2,3,4)\rangle$. Let $X=\mathbb{Z}_{(2)}^{4}\cap\{0,1\}^{4}$ be the set of $\{0,1\}$-points in $\mathbb{R}^{4}$ in layer two, i.e., each point has two $1$s and two 00s, and suppose the action is the same as in Remark 1. This action has two orbits:
$O_{G}((1,1,0,0))$ =$=\{\,(1,1,0,0),(0,1,1,0),(0,0,1,1),(1,0,0,1)\,\}\,,$
$O_{G}((1,0,1,0))$ =$=\{\,(1,0,1,0),(0,1,0,1)\,\}\,.$
The fixed space $\mathrm{Fix}_{\mathbb{Z}}(C_{4})$ contains only two $0,1$ points, namely $(1,1,1,1)$ and $(0,0,0,0)$, and neither is in $X$.
In the remainder of the paper we will need several different notions of equivalence for integral points. Potentially larger equivalence classes (based on normalizers) of core points are studied in LS2018 .
Definition 5 (Equivalence relations)
  1. 1.
    Two points $x,y\in\mathbb{Z}^{n}$ are called equivalent if there exists $g\in G$ such that $x=gy$. It follows from the group axioms that this is an equivalence relation.
  2. 2.
    Two points $x,y\in\mathbb{Z}^{n}$ are called isomorphic if there exists $g\in G$ such that $x-gy\in\mathrm{Fix}_{\mathbb{Z}}(G)$. This is an equivalence relation because $\mathrm{Fix}_{\mathbb{Z}}(G)$ is a lattice.
  3. 3.
    Two integer points $z_{1}$ and $z_{2}$ in $\mathbb{Z}^{n}$ are called co-projective if there exists an integer $k\in\mathbb{Z}$ such that $z_{1}=z_{2}+k\mathbf{1}$. Equivalently, if the group is transitive, $z_{2}$ is a translation of $z_{1}$ through the fixed space.
Each of the equivalence relations in Definition 5 has equivalence classes that are either entirely core-points or entirely non-core integer points. Along with the observation that $\{0,1\}$-vectors are core points for any permutation group (2, , Lemma 3.7), we can define a family of universal core points.
Definition 6
A point $u\in\mathbb{Z}^{n}$ is called a universal core point if it is isomorphic to a $\{0,1\}$-vector.
As a heuristic for identifying more points contained in the orbit polytopes of many (non-core) points, we consider integer points near universal core points.
Definition 7
An integer point $z$ is called an atom if there is a universal core point $u$ in the layer containing $z$ such that the distance between $z$ and $u$ is $\sqrt{2}$.
For example if $G=\langle(1,2,3,4,5)\rangle$ then the fixed space is spanned by $\mathbf{1}$ and $(2,2,2,2,1)=(1,1,1,1,0)+(1,1,1,1,1)$ is a universal core point. The point $(3,2,2,1,1)$ is an atom since $(3,2,2,1,1)-(2,2,2,2,1)=(1,0,0,-1,0)$ hence its distance to the universal core point $(2,2,2,2,1)$ is $\sqrt{2}$.
Definition 8
Suppose the cyclic group $G=\langle(g_{1},\ldots,g_{k})\rangle\leqslant S_{n}$, $g_{i}\in\{0,\ldots,n-1\}$ acts on $\mathbb{R}^{n}$ by permuting coordinates. Then coordinate $i$ is called active if there is $j\in\{1,\ldots,k\}$ such that $g_{j}=i$.
For example if $G=\langle(0,1,3,4)\rangle$ acts on $\mathbb{R}^{5}$ then $x_{0},x_{1},x_{3},x_{4}$ are active but $x_{2}$ is non-active.
3 Circulant Matrices
Circulant matrices play an important role in finding our constraints because any point $x$ in the orbit polytope of the integer point $c\in\mathbb{Z}^{n}$ under the cyclic group $C_{n}$ can be written as $x=C\lambda$, where $x\in\mathbb{R}^{n}$, $\lambda\in[0,1]^{n}$ and $C$ is the circulant matrix of $c$.
Definition 9
A circulant matrix is a matrix where each column vector is rotated one element down relative to the preceding column vector. An $n\times n$ circulant matrix $\operatorname{Cir}(c)$ takes the form
$C=\begin{bmatrix}c_{0}&c_{n-1}&\ldots&c_{2}&c_{1}\\ c_{1}&c_{0}&c_{n-1}&\ldots&c_{2}\\ \vdots&c_{1}&c_{0}&\ddots&\vdots\\ c_{n-2}&&\ddots&\ddots&c_{n-1}\\ c_{n-1}&c_{n-2}&\cdots&c_{1}&c_{0}\end{bmatrix}.$
One amazing property of circulant matrices is that the eigenvectors are always the same for all $n\times n$ circulant matrices. The eigenvalues are different for each matrix, but since we know the eigenvectors a circulant matrix can be diagonalized easily. For more detailed background on circulant matrices see Cir .
The $m$-th eigenvector $y^{m}$ for any $n\times n$ circulant matrix $\operatorname{Cir}(c)$ is given by:
$y^{m}=\frac{1}{\sqrt{n}}(1,w^{-m}_{n},\ldots,w_{n}^{-(n-1)m})^{T},$ (1)
where $w^{m}_{n}=\exp(2\pi mi/{n}).$ Suppose $c=(c_{0},\ldots{},c_{n-1})\in\mathbb{R}^{n}$ (usually for us $c$ will be an integer point in $\mathbb{Z}^{n}$). By Euler’s formula we have $\sqrt{n}y^{m}=V_{m}-iU_{m},\;\;m=0,\ldots{},n-1$, where
$V_{m}=\left(1,\cos\left(\frac{2\pi m}{n}\right),\ldots{},\cos\left(\frac{2\pi(n-1)m}{n}\right)\right),$ (2)
$U_{m}=\left(0,\sin\left(\frac{2\pi m}{n}\right),\ldots{},\sin\left(\frac{2\pi(n-1)m}{n}\right)\right).$ (3)
The eigenvalue of the $m$-th eigenvector of the circulant matrix $\operatorname{Cir}(c)$ is
$\psi_{m}$ $=\sum_{k=0}^{n-1}c_{k}w^{km}_{n}=\langle V_{m},c\rangle+i\langle U_{m},c\rangle.$
So we have
$\operatorname{Cir}(c)Y$ $=Y\Psi\;\;\;\;\Longrightarrow\;\;\;\;\;\operatorname{Cir}(c)=Y\Psi Y^{*},$
where
$Y$ $=[y^{0}\mid\ldots\mid y^{n-1}],$
is the unitary matrix composed of the eigenvectors as columns, and $\Psi$ is the diagonal matrix with diagonal elements $\psi_{0},\ldots,\psi_{n-1}$.
The inverse of a circulant matrix is circulant Book and its inverse is given by
$\operatorname{Cir}(c)^{-1}=Y\Psi^{-1}Y^{*}.$ (4)
Since $\Psi$ is a diagonal matrix its inverse is also a diagonal matrix with diagonal elements $\psi^{-1}_{m}$, $m=0,\ldots,n-1$ where
$\dfrac{1}{\psi_{m}}=\dfrac{1}{\sum_{k=0}^{n-1}c_{k}w^{km}_{n}}=\dfrac{1}{\langle V_{m},c\rangle+i\langle U_{m},c\rangle}=\dfrac{\langle V_{m},c\rangle-i\langle U_{m},c\rangle}{\langle V_{m},c\rangle^{2}+\langle U_{m},c\rangle^{2}}.$ (5)
Remark 2
Note that the length of the projection of a vector $c\in\mathbb{R}^{n}$ onto a complex vector $v=a+ib\in\mathbb{C}^{n}$ is defined as
$\|\mathrm{Proj}^{c}_{v}\|^{2}=\dfrac{\|\langle c,a\rangle-i\langle c,b\rangle\|^{2}}{\|v\|^{2}}=\dfrac{\langle c,a\rangle^{2}+\langle c,b\rangle^{2}}{\|v\|^{2}}.$ (6)
Furthermore, the term $\langle V_{m},c\rangle^{2}+\langle U_{m},c\rangle^{2}$ in (5) is the length of the projection of $c$ onto invariant subspace $y^{m}$.
Lemma 1
Let $c\in\mathbb{R}^{n}$ and suppose $\operatorname{Cir}(c)$ is invertible. Then its inverse is $\operatorname{Cir}(\widehat{T}(c))$, where $\widehat{T}(c)$ is defined as follows:
$\widehat{T}(c)=\dfrac{1}{n}\begin{bmatrix}\dfrac{1}{\langle c,\mathbf{1}\rangle}+T_{0}(c)\\ \dfrac{1}{\langle c,\mathbf{1}\rangle}+T_{1}(c)\\ \vdots\\ \dfrac{1}{\langle c,\mathbf{1}\rangle}+T_{n-1}(c)\end{bmatrix}=\dfrac{1}{n}\begin{bmatrix}\dfrac{1}{\langle c,\mathbf{1}\rangle}+\psi^{-1}_{1}+\ldots+\psi^{-1}_{n-1}\\ \dfrac{1}{\langle c,\mathbf{1}\rangle}+w_{n}^{-1}\psi^{-1}_{1}+\ldots+w^{-(n-1)}_{n}\psi^{-1}_{n-1}\\ \vdots\\ \dfrac{1}{\langle c,\mathbf{1}\rangle}+w^{-(n-1)}_{n}\psi^{-1}_{1}+..+w^{-(n-1)^{2}}_{n}\psi^{-1}_{n-1}\end{bmatrix}.$
Note that actually $\dfrac{1}{\langle c,\mathbf{1}\rangle}=\psi_{0}^{-1}.$
Proof
By (4) we have $\operatorname{Cir}(c)^{-1}=Y\Psi^{-1}Y^{*}$. Now suppose $k_{hj}$ is the $(h,j)$-th component of $Y\Psi^{-1}$. We have
$k_{hj}=Y_{hj}\Psi_{jj}^{-1}=\dfrac{1}{\sqrt{n}}w^{-(h-1)(j-1)}_{n}\Psi^{-1}_{(j-1)(j-1)}.$
So, $k^{l}$, the $l$-th row of $Y\Psi^{-1}$, is equal
$k^{l}=\dfrac{1}{\sqrt{n}}\left[\Psi_{00}^{-1},w^{-(l-1)}_{n}\Psi^{-1}_{11},w^{-2(l-1)}_{n}\Psi^{-1}_{22},\ldots,w_{n}^{-(n-1)(l-1)}\Psi^{-1}_{(n-1)(n-1)}\right]\,.$
Now since $\operatorname{Cir}(c)^{-1}$ is a circulant matrix Book it is enough to find the first column of $\operatorname{Cir}(c)^{-1}$ (the other columns can be found by permutation of the first column). Notice that the first row and column of $Y$ and $Y^{*}$ is $\frac{1}{\sqrt{n}}\mathbf{1}=\frac{1}{\sqrt{n}}(1,1,\ldots,1)$ so multiplying each row of $Y\Psi^{-1}$ with vector $\frac{1}{\sqrt{n}}\mathbf{1}$ gives us the first column of $Y\Psi^{-1}Y^{*}$, which is
$\langle k^{l},\frac{1}{\sqrt{n}}\mathbf{1}\rangle=\frac{1}{n}(\Psi^{-1}_{00}+w^{-(l-1)}_{n}\Psi^{-1}_{11}+w^{-2(l-1)}_{n}\Psi^{-1}_{22}+\ldots+w^{-(n-1)(l-1)}_{n}\Psi^{-1}_{(n-1)(n-1)}),$
where $\Psi^{-1}_{00}={1}/{\langle c,\mathbf{1}\rangle}.$
The following theorem plays an important role in this paper, and is a useful formula for computing the inverse of an invertible circulant matrix.
Theorem 3.1
For $c\in\mathbb{R}^{n}$, the entry $T_{k}(c)$ in the column vector $\widehat{T}(c)$ in Lemma 1 can be written as
$T_{k}(c)$ $=2\sum_{m=1}^{(n-1)/2}\dfrac{1}{\langle V_{m},c\rangle^{2}+\langle U_{m},c\rangle^{2}}\langle\sigma^{-k}(V_{m}),c\rangle\;\;\;\;\;\text{if n is odd},$
$T_{k}(c)$ $=2\sum_{m=1}^{(n-2)/{2}}\dfrac{1}{\langle V_{m},c\rangle^{2}+\langle U_{m},c\rangle^{2}}\langle\sigma^{-k}(V_{m}),c\rangle+\dfrac{(-1)^{k}}{\langle V_{\frac{n}{2}},c\rangle}\;\;\;\text{if n is even}.$
In particular each $T_{k}(c)$ is a real number.
Proof
As shown in Lemma 1 we have
$T_{k}(c)=w^{-k}_{n}\psi^{-1}_{1}+\ldots+w^{-(n-1)k}_{n}\psi^{-1}_{n-1}.$
Since $\psi_{l}$ and $\psi_{n-l}$ are complex conjugates of each other we have
$w^{-lk}_{n}\psi^{-1}_{l}+w^{-(n-l)k}_{n}\psi_{n-l}^{-1}=(\cos(\frac{2lk\pi}{n})-i\sin(\frac{2lk\pi}{n}))\dfrac{\langle c,V_{l}\rangle-i\langle c,U_{l}\rangle}{\langle c,V_{l}\rangle^{2}+\langle c,U_{l}\rangle^{2}}\\ +(\cos(\frac{2lk\pi}{n})+i\sin(\frac{2lk\pi}{n}))\dfrac{\langle c,V_{l}\rangle+i\langle c,U_{l}\rangle}{\langle c,V_{l}\rangle^{2}+\langle c,U_{l}\rangle^{2}}=\\ \dfrac{2\cos(\frac{2lk\pi}{n})\langle c,V_{l}\rangle-2\sin(\frac{2lk\pi}{n})\langle c,U_{l}\rangle}{\langle c,V_{l}\rangle^{2}+\langle c,U_{l}\rangle^{2}}=2\dfrac{\langle c,\sigma^{-k}(V_{l})\rangle}{\langle c,V_{l}\rangle^{2}+\langle c,U_{l}\rangle^{2}}.$
Recall that $w^{m}_{n}=\cos\left(\frac{2\pi m}{n}\right)+i\sin\left(\frac{2\pi m}{n}\right)$ so the last equality holds because $\cos A\cos B-\sin A\sin B=\cos\left(A+B\right)$ and $V_{m}$ is in terms of $\cos$ and $U_{m}$ is in terms of $\sin$ (see (2) and (3)).
If $n$ is even then for $l=\frac{n}{2}$ we have $\psi_{\frac{n}{2}}=\psi_{n-l}$ and $y^{\frac{n}{2}}=y^{n-l}$ so $w^{-\frac{n}{2}k}_{n}\psi^{-1}_{\frac{n}{2}}$ does not give a complex conjugate pair. But since the imaginary part of it is zero, we have
[…]
Lemma 2
Suppose for $c\in\mathbb{R}^{n}$ that $\operatorname{Cir}(c)$ is invertible. Then $\sum_{k=0}^{n-1}T_{k}(c)=0$.
Proof
From Lemma 1 we have
$\sum_{k=0}^{n-1}T_{k}(c)$ $=\sum_{k=0}^{n-1}\sum_{m=1}^{n-1}w_{n}^{-km}\psi_{m}^{-1}=\sum_{m=1}^{n-1}\psi_{m}^{-1}\sum_{k=0}^{n-1}w_{n}^{-km}\,.$
Since the inner sum is geometric, for $m<n$
$\sum_{k=0}^{n-1}w_{n}^{-km}$ $\qedhere\displaystyle=0\,.\qedhere$
Remark 3
Note that if $z,c\in\mathbb{Z}_{(k)}^{n},\;k\neq 0,$ and $z=\operatorname{Cir}(c)\lambda$ then
$k=\langle\mathbf{1},z\rangle=\langle\mathbf{1},\operatorname{Cir}(c)\lambda\rangle=\langle\mathbf{1},c\rangle\langle\mathbf{1},\lambda\rangle=k\langle\mathbf{1},\lambda\rangle.$
This implies that $\langle\mathbf{1},\lambda\rangle=1.$ Now suppose $\operatorname{Cir}(c)$ is invertible, so $\lambda=\operatorname{Cir}(c)^{-1}z$. To check if $z\in\mathrm{conv}(G_{c})$ or not, we only need to check if $\lambda=\operatorname{Cir}(c)^{-1}z\geqslant\mathbf{0}$. The constraint $\sum_{i=0}^{n-1}\lambda_{i}=1$ is redundant.
Let us denote by $\overline{T}(c)$ the first row of $\operatorname{Cir}(\widehat{T}(c))$ which is
$\overline{T}(c)=\dfrac{1}{n}\left[\dfrac{1}{\langle c,\mathbf{1}\rangle}+T_{0}(c),\dfrac{1}{\langle c,\mathbf{1}\rangle}+T_{n-1}(c),\ldots{},\dfrac{1}{\langle c,\mathbf{1}\rangle}+T_{1}(c)\right].$
This will simplify the notation in following sections.
3.1 New Constraints for Singular Circulant Matrices
The circulant matrix $\operatorname{Cir}(c)$ corresponding to an integer point $c\in\mathbb{Z}^{n}$ is not invertible if and only if the determinant of $\operatorname{Cir}(c)$ is zero. Since the determinant of a square matrix is equal to the product of its $n$ eigenvalues we have
$\det(\operatorname{Cir}(c))=\prod_{j=0}^{n-1}(\langle V_{j},c\rangle+i\langle U_{j},c\rangle)\,.$
Furthermore, since for $k=1,\ldots{},n-1$, $V_{n-k}+iU_{n-k}$ is the complex conjugate of $V_{k}+iU_{k}$ we have
$\det(\operatorname{Cir}(c))=\begin{cases}\langle V_{0},c\rangle\prod_{j=1}^{(n-1)/{2}}(\langle V_{j},c\rangle^{2}+\langle U_{j},c\rangle^{2})&n\text{ odd,}\\[4.73611pt] \langle V_{0},c\rangle\langle V_{\frac{n}{2}},c\rangle\prod_{j=1}^{(n-2)/{2}}(\langle V_{j},c\rangle^{2}+\langle U_{j},c\rangle^{2})&n\text{ even.}\end{cases}$
So in order to check if an ILP has an integer solution whose circulant matrix is singular, we could add the following constraints to the problem
$\langle V_{0},c\rangle\prod_{j=1}^{(n-1)/{2}}(\langle V_{j},c\rangle^{2}+\langle U_{j},c\rangle^{2})$ $=0\;\;\;\;\;\;\;\text{if}\;n\;\text{is odd},$ (7)
$\langle V_{0},c\rangle\langle V_{\frac{n}{2}},c\rangle\prod_{j=1}^{(n-2)/{2}}(\langle V_{j},c\rangle^{2}+\langle U_{j},c\rangle^{2})$ $=0\;\;\;\;\;\;\text{if}\;n\;\text{is even}.$ (8)
Equations (7) and (8) involve polynomials in $c$ of degree $n$. We can reformulate them in a way that make them easier to solve in practice by introducing a large positive constant $M$ and ${0,1}$-valued variables $r_{m}$. Let $P_{j}$ be a term in the products (7) and (8), namely, $P_{j}=\langle V_{j},c\rangle$ when $j=0$ or (for $n$ even) when $j=\frac{n}{2}$. Otherwise, $P_{j}=\langle V_{j},c\rangle^{2}+\langle U_{j},c\rangle^{2}\geqslant 0$. For $j=0,\ldots{},\lceil({n-1})/{2}\rceil$, add constraints
$P_{j}$ $\leqslant r_{j}P_{j}$ $j\not\in\{0,\frac{n}{2}\}\cap\mathbb{Z},$ (9)
[…]
Finally add the following constraint to force at least one $r_{j}$ to 00:
$\sum_{j=0}^{\left\lceil(n-1)/{2}\right\rceil}r_{j}\leqslant\left\lceil\frac{n-1}{2}\right\rceil\,.$ (10)
Note that constraints (9) and (10) forces at least one of the $P_{j}$ to be zero and so the determinant will be zero. For maximization problems the constant $M$ can be chosen as the absolute value of the objective value of the LP relaxation (assuming the objective function $f(x)=\langle\mathbf{1},x\rangle$).
Remark 4
We can formulate (7) and (8) in different ways. For example we can make $\left\lceil\frac{n-1}{2}\right\rceil$ subproblems by adding each $P_{m}$ separately. The corresponding constraints in each subproblem are linear. For example in the $h$-th subproblem, $h\in\{1,\ldots,\left\lceil\frac{n-1}{2}\right\rceil\}$, we have
$P_{h}=\langle V_{h},c\rangle^{2}+\langle U_{h},c\rangle^{2}=0,$
which can be simplified as
$\langle V_{h},c\rangle=0\text{ and }\langle w_{h},c\rangle=0.$
The weakness of this formulation is that adding these constraints does not simplify the problem enough because in each step we are searching in $n-2$ dimensions which for large $n$ is not a sufficient dimension reduction.
3.2 New Constraints for Non-Singular Circulant Matrices
In this section we develop some new constraints to find an outer approximation of core points. These new constraints depend only on the symmetry group and not the ILP.
Suppose $c$ is an arbitrary integer point in ($\mathbb{Z}^{n}_{(k)},\;k\neq 0,$ and $G$ is the cyclic group with order $n$ which acts on coordinates as usual. By Remark 3 a point $z\in\mathbb{Z}^{n}_{(k)}$ is in $\mathrm{conv}(G_{c})$ if and only if there exist $\lambda=(\lambda_{0},\ldots,\lambda_{n-1})$ such that
$\SwapAboveDisplaySkip z$ $=\operatorname{Cir}(c)\lambda,\qquad\lambda_{i}\geqslant 0.$
If $\operatorname{Cir}(c)$ is invertible we have
$\lambda$ $=\operatorname{Cir}(c)^{-1}z,$
and by Lemma 1
$\lambda_{j}$ $=\langle(\sigma^{j}(\overline{T}(c)),z\rangle=\langle\overline{T}(c),\sigma^{-j}(z)\rangle\,.$ (11)
The orbit polytope of a core point is lattice-free. Still assuming that $\operatorname{Cir}(c)$ is invertible, we conclude that $c\in\mathbb{Z}^{n}_{(k)}$ is a core point if and only if for each $z\in\mathbb{Z}^{n}_{(k)}$ which is not in the $G$-orbit of $c$,
$\min_{j}\;\langle\overline{T}(c),\sigma^{-j}(z)\rangle<0,$ (12)
Since $c$ and $z$ lie in the same layer $\langle c,\mathbf{1}\rangle=k$, $\lambda_{j}<0$ is equivalent to (interpreting indices modulo $n$)
$z_{n+j-0}T_{0}(c)+z_{n+j-1}T_{1}(c)+\ldots+z_{n+j-(n-1)}T_{n-1}(c)+1<0.$
Informally, we may say that for at least one permutation of subscripts, we have
$z_{0}T_{0}(c)+z_{1}T_{1}(c)+\ldots+z_{n-1}T_{n-1}(c)+1<0.$ (13)
It should be mentioned that, for almost all $z\in\mathbb{Z}^{n}$ the constraint $\langle z,\overline{T}(c)\rangle<0$ is nonlinear and non-convex with respect to $c$. Considering a single $z$ in isolation, we can assume $\lambda_{0}$ is negative, because if for a core point $c$, $\lambda_{0}$ is not negative, there is a permutation of $c$ that has negative $\lambda_{0}$.
Define
$H(z)=z_{0}T_{0}(c)+z_{1}T_{1}(c)+\ldots+z_{n-1}T_{n-1}(c)+1,$
so that $H(z)=0$ is the equation of a hyperplane in $\mathbb{R}^{n}$ with non-zero normal vector $(T_{0}(c),T_{1}(c),\ldots,T_{n-1}(c))$ which is perpendicular to the fixed space (Lemma 2). (If all $T_{j}(c)=0$, then $\widehat{T}(c)$ in Lemma 1 is a multiple of $\mathbf{1}$. But this makes $\operatorname{Cir}(c)^{-1}$ non-invertible). More precisely, $c\in\mathbb{Z}^{n}_{(k)}$ is a core point if for all integer points $z\in\mathbb{Z}^{n}_{(k)}$ there is an index $j\in[n]$ such that $H(\sigma^{j}(z))<0$.
Suppose $Q$ is the orbit polytope of a non-core point and $R$ is the orbit polytope of a universal core point in the same layer. Intuitively, we expect that many integer points whose orbit polytope contains $Q$, also have orbit polytope containing $R$. The idea for making new constraints is to remove from the feasible region all integer points whose orbit polytopes contain atoms or universal core points. This process can be done by searching layer by layer in the feasible region.
Lemma 3
For two co-projective integer points $c$,$c^{\prime}\in\mathbb{Z}^{n}$ we have
$T_{j}(c)=T_{j}(c^{\prime})\text{ for all }\;j=0,\ldots,n-1.$
Moreover, the constraints (12) are invariant under translation in the fixed space. That is, if $z$ and $z^{\prime}$ are two co-projective integer points in the same non-zero layers as $c$ and $c^{\prime}$ respectively, we have
$\langle z,\overline{T}(c)\rangle=\langle z^{\prime},\overline{T}(c^{\prime})\rangle.$
Proof
Recall from Definition 5 that $c^{\prime}=c+k\mathbf{1}$ for some $k\in\mathbb{Z}$. Since $\mathbf{1}$ is orthogonal to the other eigenvectors $y^{m}$, $\;m=1,\ldots,n-1$, we have
$\psi_{m}$ $=\sqrt{n}\langle c,y^{m}\rangle=\sqrt{n}\langle c+k\mathbf{1},y^{m}\rangle=\sqrt{n}\langle c^{\prime},y^{m}\rangle$
$=\psi^{\prime}_{m}\;\;\;\;\forall m=1,\ldots,n-1$
$\psi_{0}$ $=\langle c,\mathbf{1}\rangle,\;\;\;\psi_{0}^{\prime}=\langle c^{\prime},\mathbf{1}\rangle=\langle c,\mathbf{1}\rangle+\langle k\mathbf{1},\mathbf{1}\rangle\,.$
Furthermore $\operatorname{Cir}(c)$ and $\operatorname{Cir}(c^{\prime})$ have different inverses. But by Lemma 1 $T_{j}(c)=\sum_{m=1}^{n-1}w^{-jm}_{n}\psi^{-1}_{m}$. Since $\psi_{m}=\psi_{m}^{\prime}$ we get $T_{j}(c)=T_{j}(c^{\prime}),\;j=0,\ldots,n-1.$ Now let $z\in\mathbb{Z}^{n}$ and $z^{\prime}=z+k\mathbf{1}$ be in the same layers as $c$ and $c^{\prime}$ respectively. With the help of Lemma 2 we have
$n\langle z,\overline{T}(c)\rangle$ $=z_{0}T_{0}(c)+z_{1}T_{n-1}(c)+\dots+z_{n-1}T_{1}(c)+1$
$=z_{0}T_{0}(c)+\dots+z_{n-1}T_{1}(c)+k(T_{0}(c)+\dots+T_{n-1}(c))+1$
$=(z_{0}+k)T_{0}(c)+(z_{1}+k)T_{n-1}(c)+\dots+(z_{n-1}+k)T_{1}(c)+1$
$\qedhere\displaystyle=z_{0}^{\prime}T_{0}(c)+z_{1}^{\prime}T_{n-1}(c)+\dots+z^{\prime}_{n-1}T_{1}(c)+1=n\langle z^{\prime},\overline{T}(c^{\prime})\rangle\,.\qedhere$ (14)
Figure 1: Illustrating the proof of Lemma 4.
The main idea of our algorithm for the transitive case (all variables active) is as follows.
Lemma 4
Let $P$ be an integer linear maximization problem with $n$ variables, objective function $f(x)=\langle\mathbf{1},x\rangle$ and transitive symmetry group $G$. Let the LP relaxation of $P$ have feasible region $R$ and optimal solution $x^{*}$. Let $P_{j}$ be the integer feasibility problem obtained by intersecting $R$ with layer $j$. Let $l=\lfloor f(x^{*})\rfloor$. Define $F=\{l-n<j\leqslant l\mid P_{j}\text{ is feasible}\}$. If $F$ is empty, then $P$ is infeasible; otherwise the optimal solution is in layer $k=\max(F)$.
Proof
In the following discussion, we use above and below to refer to the natural ordering of layers by index. Since we assume $G$ acts by permuting coordinates, there is no real loss of generality in fixing the objective function to $f(x)=\langle\mathbf{1},x\rangle$. As we already observed, since $G$ is transitive $\mathrm{Fix}_{\mathbb{R}}(G)=\mathrm{Span}(\mathbf{1})$. It is clear that layer $l$ is the highest layer which could intersect $R$; furthermore if $F$ is non-empty the optimal solution of $P$ must lie in the highest layer with non-empty intersection with $R$. Suppose $F$ is empty, so no layer above $l-n$ is feasible.
Define the orbit-barycenter $b(x)$ for $x\in\mathbb{R}^{n}$ as
$b(x)=\dfrac{1}{|G|}\sum_{g\in G}gx.$
Note that this point lies in $\mathrm{Fix}_{\mathbb{R}}(G)$ and $\langle\mathbf{1},x\rangle=\langle\mathbf{1},b(x)\rangle$; furthermore if $x\in R$ then $b(x)\in R$. This allows us to assume w.l.o.g.  that $x^{*}\in\mathrm{Fix}_{\mathbb{R}}(G)$.
There exists integer $k$ satisfying $l-n<l^{\prime}=kn\leqslant l$. The integer point $k\mathbf{1}$ is the unique intersection of layer $l^{\prime}$ and $\mathrm{Fix}_{\mathbb{Z}}(G)$, so since $P_{l^{\prime}}$ is infeasible, $R$ does not intersect layer $l^{\prime}$. Convexity then tells us that no layer below $l^{\prime}$ intersects $R$ (as the segment of $\mathrm{Fix}_{\mathbb{R}}(G)$ between such a layer and $x^{*}$ goes through $k\mathbf{1}$).
Note that in order to check if an integer point $c\in\mathbb{Z}^{n}_{(k)}$ is a core point or not it is impossible to add constraints of the form of (12) for all $z\in\mathbb{Z}^{n}_{(k)}$ because there are infinitely many integer points in each layer. But we can add constraints for a finite subset of integer points in ($\mathbb{Z}^{n}_{(k)}$. As mentioned earlier, it makes sense to choose atom points and universal core points since they are the closest integer points to the barycenter. For this finite set we have the following definition.
Definition 10
In each layer $k$, any choice of set of atoms and universal core points for making inequalities (12) is called an essential set of the layer $k$ and is denoted by $E^{k}$.
Note that since vectors in $\{0,1\}^{n}$ are core points, a non-empty choice of essential set always exists.
Definition 11
We say that layers $l$ and $l^{\prime}$ in $\mathbb{Z}^{n}$ are congruent if $l\equiv l^{\prime}$ $\pmod{n}$.
Lemma 2 plays an important role in defining the specific essential set for any layer.
Remark 5
We can translate each integer point $z\in E^{k}$ through the fixed space to get an integer point with entries in $\{-1,-2,0,1,2\}$ and use inequalities (12), which can be written as
$01+\langle\sigma^{-j}(z),(T_{0}(c),T_{n-1}(c),\ldots{},T_{1}(c))\rangle<$ 0 . (15)
Since $T_{0}(c)+T_{1}(c)+\ldots+T_{n-1}(c)=0$, the inequality (15) holds for both or neither co-projective points $z,z^{\prime}$ in congruent layers.
For example, for an ILP in $\mathbb{R}^{6}$, we can define an essential set in layer $21$ by:
$E^{21}=\{(4,4,4,3,3,3),(4,3,4,4,3,3),(4,3,4,3,4,3),(5,3,4,3,3,3)\}.$
The corresponding constraint for $z=(4,4,4,3,3,3)$, $j=0$ is
$01+4T_{0}(c)+4T_{5}(c)+4T_{4}(c)+3T_{3}(c)+3T_{2}(c)+3T_{1}(c)<$ 0 . (16)
Since by Lemma 2 we have $T_{0}(c)+T_{1}(c)+T_{2}(c)+T_{3}(c)+T_{4}(c)+T_{5}(c)=0$, inequality (16) can be written as
$01+T_{0}(c)+T_{5}(c)+T_{4}(c)<$ 0 .
Furthermore, we can define the essential set as follows and use inequality (15) for making new constraints.
$E^{3}=\{(1,1,1,0,0,0),(1,0,1,1,0,0),(1,0,1,0,1,0),(2,0,1,0,0,0)\}.$
If we consider layers $1,\ldots{},n$ as representatives for all congruence classes, we can define the essential set for layers $1,\ldots{},n$ and use universal or atom points with entries $\{-2,-1,0,1,2\}$.
Definition 12
The essential set in any layers $1,\ldots{},n$ is called the projected essential set and it is denoted by $\widehat{E}^{k}$ for $k=1,\ldots{},n$.
The idea is that first we search for integer points whose circulant matrix is singular. Next we consider the case where the circulant matrix is non-singular: by adding new constraints in each layer $k$ we search for an integer point $c$ whose orbit polytope does not contain the integer points of the essential set $E^{k}$. Since $\operatorname{Cir}(c)$ is non-singular in this step, we add the following constraints (cf. Theorem 3.1 and (7), (8)):
$\langle V_{m},c\rangle^{2}+\langle U_{m},c\rangle^{2}>0\;\;\;\;\;\;\forall m=1,\ldots,\lceil(n-1)/2\rceil.$ (17)
We will present variations on this idea for different kinds of symmetry groups. In all of these variations there are three different types of subproblems as follows:
  1. ($Q_{1}$)
    Add constraints (12) (or (13)) and (17) for each point in the projected essential set.
  2. ($Q_{2}$)
    Add constraints (9) and (10) for the case of singular circulant matrices.
  3. ($Q_{3}$)
    Check the feasibility of integer points in the essential sets. For Algorithm 1 they can be tested individually, for Algorithms 2 and 3 see Remark 7.
Algorithm 1 follows the proof of Lemma 4 for the case where all variables are active.
Algorithm 1 Maximize with all variables active
Maximization ILP $P$with bounded relaxation, transitive symmetry group, and objective function $f$. Optimal objective value, or $\infty$- ∞if infeasible.
1:Construct and solve a subproblem $P_{0}$ of type $Q_{2}$ by adding constraints (9) and (10) to $P$.
2:If $P_{0}$ has integer optimal solution $\widehat{z}_{1}$, set $f_{1}=f(\widehat{z}_{1})$ otherwise set $f_{1}=-\infty$.
[…]
Here are some observations related to Algorithm 1:
  1. 1.
    There might be some non-core integer points which satisfy constraints (12) for the given essential set.
  2. 2.
    By Lemma 2, in each layer $L$ we need at most $\left\lceil(n-1)/2\right\rceil$ terms $T_{k},$ to make constraints (12). For example if $n=6$, in layer $5$ for universal core point $(1,1,1,1,1,0)$ we can use $01-T_{5}<$ 0 rather than $01+T_{0}+T_{1}+T_{2}+T_{3}+T_{4}<$ 0 .
The subproblems in steps 1 and 6 of Algorithm 1 are NP-hard, so there are no known worst case efficient algorithms. On the other hand, the cost of generating constraints, solving the LPs, and testing points for feasibility is all negligible. Actually, solving a symmetric LP can be done faster since it is enough to search through the fixed space (cf. (BHJ2013, , Theorem 1)). Since adding new constraints in steps 1 and 6 creates a nonlinear and non-convex mixed integer program, we should use a solver that can handle these constraints. In theory solving a nonlinear integer program is harder than the linear one, but in practice we see that (cf. Section 7), decreasing the feasible region by adding our constraints make it worthwhile to do so. It should also be mentioned that there is a trade off between the number of integer points in the essential set and running time of solving an ILP. Adding too many constraints might make the problem hard to solve. Some universal or atom points might be more effective for a specific problem. However, in our experiments there is no significant difference between different selected integer points in the essential sets. We summarize this discussion as follows.
Remark 6
  1. 1.
    The main cost of Algorithm 1 is in steps 1 and 6 where nonlinear integer programs must be solved.
  2. 2.
    The choice of essential set (as long as not too large) seems not to have a large impact.
3.3 Example: solving a symmetric ILP with Algorithm 1
In this section, we illustrate using Algorithm 1 to solve a small ILP. Each of the subproblems discussed here is solved on commodity hardware in at most two seconds using Knitro knitro . The integer linear program P1 with cyclic symmetry group $C_{5}$ is defined LS2018 as follows.
minimize ${\text{minimize }x_{0}+x_{1}+x_{2}+x_{3}+x_{4}}$
subj. to
$515161x_{0}+18376x_{1}-503804x_{2}-329744x_{3}+300011x_{4}$ $\leqslant 59$
$300011x_{0}+515161x_{1}+18376x_{2}-503804x_{3}-329744x_{4}$ $\leqslant 59$
$-329744x_{0}+300011x_{1}+515161x_{2}+18376x_{3}-503804x_{4}$ $\leqslant 59$
$-503804x_{0}-329744x_{1}+300011x_{2}+515161x_{3}+18376x_{4}$ $\leqslant 59$
$18376x_{0}-503804x_{1}-329744x_{2}+300011x_{3}+515161x_{4}$ $\leqslant 59$
${x_{0}+x_{1}+x_{2}+x_{3}+x_{4}=1}$
${x_{0},x_{1},x_{2},x_{3},x_{4}\in\mathbb{Z}}$
[…]
$V_{1}$ $=(1,\cos(\frac{-2\pi}{5}),\cos(\frac{-4\pi}{5}),\cos(\frac{-6\pi}{5}),\cos(\frac{-8\pi}{5}))\,,$
$V_{2}$ $=(1,\cos(\frac{-4\pi}{5}),\cos(\frac{-8\pi}{5}),\cos(\frac{-12\pi}{5}),\cos(\frac{-16\pi}{5}))\,,$
$U_{1}$ $=(0,\sin(\frac{-2\pi}{5}),\sin(\frac{-4\pi}{5}),\sin(\frac{-6\pi}{5}),\sin(\frac{-8\pi}{5}))\,,$
$U_{2}$ $=(0,\sin(\frac{-4\pi}{5}),\sin(\frac{-8\pi}{5}),\sin(\frac{-12\pi}{5}),\sin(\frac{-16\pi}{5}))\,.$
Now if $c=(c_{0},c_{1},c_{2},c_{3},c_{4})$ is an integer solution of the above ILP then the corresponding $T_{i}$ in Theorem 3.1 are as below:
$T_{0}$ $=\dfrac{\langle V_{1},c\rangle}{\langle V_{1},c\rangle^{2}+\langle U_{1},c\rangle^{2}}+\dfrac{\langle V_{2},c\rangle}{\langle V_{2},c\rangle^{2}+\langle U_{2},c\rangle^{2}}\,,$
$T_{1}$ $=\dfrac{\langle\sigma^{-1}(V_{1}),c\rangle}{\langle V_{1},c\rangle^{2}+\langle U_{1},c\rangle^{2}}+\dfrac{\langle\sigma^{-1}(V_{2}),c\rangle}{\langle V_{2},c\rangle^{2}+\langle U_{2},c\rangle^{2}}\,,$
$T_{2}$ $=\dfrac{\langle\sigma^{-2}(V_{1}),c\rangle}{\langle V_{1},c\rangle^{2}+\langle U_{1},c\rangle^{2}}+\dfrac{\langle\sigma^{-2}(V_{2}),c\rangle}{\langle V_{2},c\rangle^{2}+\langle U_{2},c\rangle^{2}}\,,$
$T_{3}$ $=\dfrac{\langle\sigma^{-3}(V_{1}),c\rangle}{\langle V_{1},c\rangle^{2}+\langle U_{1},c\rangle^{2}}+\dfrac{\langle\sigma^{-3}(V_{2}),c\rangle}{\langle V_{2},c\rangle^{2}+\langle U_{2},c\rangle^{2}}\,,$
$T_{4}$ $=\dfrac{\langle\sigma^{-4}(V_{1}),c\rangle}{\langle V_{1},c\rangle^{2}+\langle U_{1},c\rangle^{2}}+\dfrac{\langle\sigma^{-4}(V_{2}),c\rangle}{\langle V_{2},c\rangle^{2}+\langle U_{2},c\rangle^{2}}\,.$
Since all variables are active we can use Algorithm 1 to solve the problem. Note that since the problem is in layer 2, we do not need to solve an LP relaxation problem to determine layers.
First we choose the essential set in layer 2 as $E^{2}=\{(1,1,0,0,0)\}$. By adding the following nonlinear constraint,
$1+T_{0}(c)+T_{1}(c)<0,$
we quickly (via Knitro) see the subproblem is infeasible. Checking the feasibility of (1,1,0,0,0) is trivial. By adding the following constraints (see (9) and (10)) we check if there is an integer solution that has a singular circulant matrix.
$\langle V_{1},c\rangle^{2}+\langle U_{1},c\rangle^{2}$ $\leqslant r_{1}(\langle V_{1},c\rangle^{2}+\langle U_{1},c\rangle^{2})$
$\langle V_{2},c\rangle^{2}+\langle U_{2},c\rangle^{2}$ $\leqslant r_{2}(\langle V_{2},c\rangle^{2}+\langle U_{2},c\rangle^{2})$
${-2r_{3}\leqslant c_{0}+c_{1}+c_{2}+c_{3}+c_{4}\leqslant 2r_{3}}$
${r_{1}+r_{2}+r_{3}\leqslant 2}$
The final subproblem was also solved quickly by Knitro, and no integer solution was found.
4 Partial-Circulant Matrices
In the previous section we assumed that the order of the cyclic permutation group of an ILP was equal to the dimension of the problem. In this section, we generalize the algorithm of the previous section for some ILP where not all variables are active.
For the remainder of this section, we denote by $x(1)$ the coordinates of $x$ that are active in the cyclic group $C_{k}$. We extend this notation below to direct products of several cyclic groups. Without loss of generality, assume $x(1)$ is the first $k<n$ coordinates of $x$. Notice that in this case the group action is not transitive. Indeed, the dimension of the fixed space is $n-k+1$ and it is spanned by the orthogonal vectors $e_{0}+\cdots+e_{k-1},e_{k},\ldots,e_{n-1}$ (using the usual basis vectors). Moreover, invariant subspaces of the action of $C_{k}$ on $\mathbb{R}^{k}$ embed naturally into invariant subspaces for the action of $C_{k}$ on $\mathbb{R}^{n}$. The following definition is a generalization of the circulant matrix that will be useful in generalizing the constraints of the previous section.
Definition 13
For $c\in\mathbb{R}^{n}$, an $n\times k$ Partial-Circulant Matrix, $\mathrm{PCir}(c)$, is a matrix where the first k rows are $\operatorname{Cir}(c(1))$ and the remaining rows are $\mathbf{1}^{T}$ scaled by the last $n-k$ elements of $c$.
Example 1
Let $c=(1,2,3,4,5)\in\mathbb{Z}^{5}$. Then the $5\times 3$ partial circulant matrix of $c$ takes the following form
$\begin{bmatrix}1&3&2\\ 2&1&3\\ 3&2&1\\ 4&4&4\\ 5&5&5\end{bmatrix}.$
Let $c\in\mathbb{Z}^{n}$ and suppose the circulant part of $\mathrm{PCir}(c)$ is invertible. Then the orbit polytope of $c$ has dimension $k-1$ and any point $x\in\mathbb{R}^{n}$ in $\mathrm{conv}(G_{c})$ can be written as
$\mathrm{PCir}(c)_{n\times k}\lambda=x\;\;\;\text{for some}\;\;\lambda\in\mathbb{R}^{k}.$ (18)
The rank of this partial-circulant matrix is $k$ and so the solution $\lambda$ of the system (18) can be determined by the circulant part.
Suppose the permutation $g\in S_{n}$ factors as the product $g=h_{1}\cdot h_{2}\cdots h_{d}$ of disjoint cycles $h_{j}$. Let $X_{1},\ldots,X_{d}$ be the canonically associated subspaces of $\mathbb{R}^{n}$. Thus the cyclic group $H_{j}=\langle h_{j}\rangle$ permutes standard basis vectors for $X_{j}$. Furthermore, each integer point $z\in\mathbb{Z}^{n}$ has a unique decomposition $z=\oplus_{j=1}^{d}z(j)$ with $z(j)\in X_{j}\cap\mathbb{Z}^{n}$. Keeping this notation, we have that the next lemma follows from the definition of a convex combination.
Lemma 5
Suppose $G=\langle g\rangle\leqslant S_{n}$, where $g=h_{1}\cdot h_{2}\cdots h_{d}$, is a product of disjoint cycles. Let $c\in\mathbb{Z}^{n}$. If the integer point $z=\oplus_{j=1}^{d}z(j)$ is in $\mathrm{conv}(G_{c})$, then for $j\in 1,\ldots,d$, we have $z(j)$ in the orbit polytope of $c(j)$ under $H_{j}$.
In this section we are only need the restricted case of $d=1$ of Lemma 5: if the orbit polytope of $c(1)=(c_{0},\ldots,c_{k-1})$ in $\mathbb{Z}^{k}$ is lattice-free then $c=(c_{0},\ldots,c_{n-1})$ is a core point in $\mathbb{Z}^{n}$. Moreover, since the sum of weights is 1 in a convex combination, if $z\in\mathbb{Z}^{n}$ is in $\mathrm{conv}(G_{c})$ then $z_{k+1}=c_{k+1}$, …, $z_{n}=c_{n}$. Furthermore, applying inequality (12), if $z$ does not lie in $\mathrm{conv}(G_{c})$, then for at least one cyclic permutation of $z(1)$ the following constraint must be satisfied
$z_{0}T_{0}(c)+z_{1}T_{k-1}(c)+\ldots+z_{k-1}T_{1}(c)+1<0.$ (19)
Recall that the above constraint is nonlinear and non-convex with respect to $c$.
For a given symmetric (maximization) ILP where not all variables are active in the cycle, let the optimal solution of its relaxation be $x^{*}$. In this case the symmetry group does not act transitively. So, searching in $k$ layers is not sufficient and the optimal objective value can be in any layer less than or equal to $\lfloor f(x^{*})\rfloor$. Nonetheless we can define a finite set of subproblems where we can apply the constraints developed in Section 3. In the following discussion, we use sub-layer $j$ to refer to the set of integer points whose first $k$ coordinates sum to $j$.
Lemma 6
Given an ILP $P$ with a cyclic symmetry group $G=\langle(1,2,\ldots{},k)\rangle\leqslant S_{n}$ acting on the first $k$ coordinates,
  1. 1.
    $P$ can be solved by solving $k$ subproblems created by adding the constraint
    $\sum_{j=0}^{k-1}x_{j}=i\pmod{k}\quad\text{for }i\in\{1\ldots k\}\,.$ (20)
  2. 2.
    We can use the constraints (12) in each subproblem.
  3. 3.
    Each of the subproblems can be modelled by adding a single integer variable and a single integer linear constraint to $P$.
Proof
Let $L=l+l^{*}$ be the layer of an integer point $z\in\mathbb{Z}^{n}$ in $P$ where $l$ is the layer of active variables and $l^{*}$ is the layer of non-active variables. The point $z$ must fall into one of the $k$ congruence classes defined by (20), which establishes the first point.
By Lemma 3 all congruent sub-layers $l$ have the same constraints for co-projective integer points (see Definition 5 and Definition 11). Indeed congruence is not needed here, only that co-projective points form equivalence classes for (12).
Although the constraint (20) is nonlinear, it is easy to model using integer-linear constraints. Consider the congruence relation between layers (see Definition 11). Let $i=1,\ldots,k$ be the representative of each class. Let $q_{i}$, be an integer solver variable. Constraint (20) can be reformulated as
$\qedhere\sum_{j=0}^{k-1}x_{j}=q_{i}k+i\,.\qedhere$
Algorithm 1 in the previous section can be modified in this case as Algorithm 2 below. The difference is we should check sub-layers $l_{i}=q_{i}k+i$ where $q_{1},\ldots{},q_{k}$ are some new integer variables (rather than a fixed layer). For example, if $k=3$, then all layers in $\mathbb{Z}^{3}$ can be classified as
$l_{1}=3q_{1}+1,\;l_{2}=3q_{2}+2,\text{ or }l_{3}=3q_{3}+3.$
Note that since the algorithm for this case is not searching layer by layer, checking the feasibility of integer points in the essential set is different. In other words, we know the layers of points we have (since by Remark 5 the integer points in each projected essential set are in the layers $1,\ldots{},k$) but they are only representatives for (many) pre-images, and we don’t know the layers of the pre-images. If $z(1)=\sigma^{i}(c(1))$ then the solution of $\operatorname{Cir}(c(1))\lambda=z(1)$ is $e_{a}$, that is:
$\lambda_{a}=1,\;\;\;\;\lambda_{i}=0\;\;\;\forall i=0,\ldots{},k-1,\;i\neq a,$
where $a\in\{0,\ldots,k-1\}$. Furthermore, by Theorem 3.1, to check if at least one of the pre-images of $z(1)$ in the projected essential set is a feasible point or not, for $a=0$, we can add the following constraints:
$\langle\sigma^{0}(z(1)),\overline{T}(c(1))\rangle$ $=1,$
$\langle\sigma^{1}(z(1)),\overline{T}(c(1))\rangle$ $=0,$
$\vdots$
$\langle\sigma^{k-1}(z(1)),\overline{T}(c(1))\rangle$ $=0.$
Remark 7
Since the coordinates of integer points in the projected essential set are 0,1,-1,2,-2, another way to check the feasibility of all translates along the fixed space (which uses only linear constraints with small coefficients) is to chose one of the coordinates $c_{i}$ as a base and write other coordinates with respect to that coordinate.
For example if $z=(1,-2,0,0,0,0,0)\in\widehat{E}^{1}$, then by choosing $c_{2}$ as a base we can add the following constraints to check the feasibility of all integer points $(1,-2,0,0,0,0,0)+t(1,1,1,1,1,1,1),t\in\mathbb{Z},$ in the sub-layers $17t+$ 1 :
$c_{0}$ $=c_{2}+1,$
$c_{1}$ $=c_{2}-2,$
$c_{3}$ $=c_{2},$
$c_{4}$ $=c_{2},$
$c_{5}$ $=c_{2},$
$c_{6}$ $=c_{2}\,.$
Algorithm 2 generalizes Algorithm 1 to the case where $k<n$ coordinates are active in the symmetry group. For conciseness, we write $\operatorname{OPT}_{\infty}(P)$ for the optimal objective value of maximization problem $P$. If $P$ is infeasible, $\operatorname{OPT}_{\infty}(P)$ is defined as $\infty$- ∞. In practice this can be implemented by a Boolean flag, and does not require any specialized arithmetic.
Algorithm 2 Maximize with first $k$ variables active.
Maximization ILP $P$with bounded relaxation, cyclic symmetry group acting on the first $k$coordinates, and objective function $f$. Optimum objective value, or $\infty$- ∞if $P$is infeasible.
[…]
2:Construct a subproblem $L_{i}$ by adding constraint $\sum_{j=0}^{k-1}x_{j}=q_{i}k+i$ where $q_{i}$ is a new integer variable.
[…]
4:Construct and solve subproblem $P^{j}_{i}$ of type $Q_{3}$ by adding constraints described in Remark 7 to $L_{i}$ for point in $z^{j}\in\widehat{E}^{i}$
[…]
6:Construct a subproblem $P_{i}$ of type $Q_{1}$ by adding to $L_{i}$ constraint (17) for all $m=0,\ldots,\left\lceil(n-1)/{2}\right\rceil$ and constraints (12) for each $z^{j}\in\widehat{E}^{i}$.
7:$f^{*}\leftarrow\max\{f^{*},\operatorname{OPT}_{\infty}(P_{i})\}.$EndFor
[…]
As discussed in Remark 6, giving precise complexity bounds for Algorithm 2 is difficult. But it is worthwhile to compare the complexity of Algorithm 1 and Algorithm 2. In step 1 of of both algorithms the same constraints are used but the difference is that, in the second algorithm there are some non-active coordinates. So the subproblem in this case is harder to solve since constraints are effective on a small dimension of the problem. The type of constraints in step 6 of Algorithm 2 and step 6 of Algorithm 1 is the same but there are two differences: 1) less coordinates are again active in the second algorithm, and 2) in the first algorithm in each subproblem is in a single layer, but in the second algorithm in each subproblem is solved over an equivalence class of sub-layers. Both of these factors make the type $Q_{1}$ subproblems harder to solve in Algorithm 2. The other important fact is that, in the first algorithm since we are searching layer by layer, the first integer solution (if there is any) is an optimal solution. But in the second algorithm since all coordinates are not active if an integer solution is found it might not be optimal. On the other hand, when using Algorithm 2 for feasibility problems the subproblems can also be solved as feasibility problems, using local solvers if desired. Finally, checking the feasibility of universal and atom core points in these two algorithms is done differently. In the first algorithm in each subproblem, it is just testing an integer point against a set of linear inequalities. But in the second algorithm since not all coordinates are active, in each subproblem an ILP should be solved after adding constraints described in Remark 7.
5 New Constraints for Direct Products of Cyclic Groups
Lemma 6 shows that Algorithm 2 can be used for a cyclic subgroup of the symmetric group of an ILP. In some cases the cyclic subgroup is small and Algorithms 1 and 2 are not very practical. In this section we generalize Algorithm 2 for direct products of cyclic groups. Like Algorithm 2, the algorithm in this section also does not search layer by layer in the feasible region. Instead, we search in all equivalence classes of sub-layers.
Recall that the Cartesian product of the $d$ sets $X_{1},\ldots,X_{d}$ is
$X_{1}\times\ldots{}\times X_{d}=\prod_{i=1}^{d}X_{i}=\{(x_{1},\ldots{},x_{d}):\;x_{i}\in X_{i}\;\text{for every }i\in\{1,\ldots{},d\}\}.$
Definition 14
Let $G_{i},i=1,\ldots{},d$ be some finite groups. The direct product $G_{1}\times\ldots{}\times G_{d}$ is defined as follows.
  1. 1.
    The underlying set is $G_{1}\times\ldots{}\times G_{d}$.
  2. 2.
    Multiplication is defined coordinate-wise:
    $(g_{1},\ldots{},g_{d})\cdot(g_{1}^{\prime},\ldots{},g_{d}^{\prime})=(g_{1}.g^{\prime}_{1},\ldots{},g_{d}.g_{d}^{\prime}).$
  3. 3.
    The identity element of this group is defined as $(e_{1},\ldots,e_{d})$ where $e_{j}$ is the identity element of $G_{j}$ for $j=1,\cdots,d.$
It is routine to check that with the above operation, $\prod_{i=1}^{d}G_{i}$ is a group. Consider a product $g=h_{1}\cdot h_{2}\cdots h_{d}$ of disjoint cycles in $S_{n}$, and let $H_{i}=\langle h_{i}\rangle$.
Lemma 7
The direct product $G=H_{1}\times\cdots\times H_{d}$ is isomorphic to $G^{\prime}=\langle h_{1},\ldots,h_{d}\rangle$.
Proof
Let $g_{i}=h_{i}^{t_{i}}$ denote a general element of the cyclic group $H_{i}$. Consider the following map between $G$ and $G^{\prime}$
$\phi:G$ $\rightarrow G^{\prime}$
$(g_{1},\ldots{},g_{d})$ $\rightarrow g_{1}g_{2}\cdots g_{d},$
It is straightforward to check $\phi$ is a bijection since the cycles are disjoint ($g_{i}$’s commute since the $h_{i}$’s are disjoint). Also we have
$\phi((g_{1},\ldots{},g_{d})\cdot(g_{1}^{\prime},\ldots{},g_{d}^{\prime}))=\phi(g_{1}g^{\prime}_{1},\ldots{},g_{d}g_{d}^{\prime})=g_{1}g^{\prime}_{1}\ldots{}g_{d}g_{d}^{\prime}=$
$g_{1}\ldots{}g_{d}g_{1}^{\prime}\ldots{}g_{d}^{\prime}=\phi(g_{1},\ldots{},g_{d})\phi(g_{1}^{\prime},\ldots{},g_{d}^{\prime}).$
So $\phi$ is a group isomorphism. Note that the second to last equality holds since the cycles are disjoint.
Example 2
The permutation group $G=\langle(1,2,3,4),(5,6,7)\rangle$ on the set $\{1,2,3,4,5,6,7\}$ is isomorphic to the direct product of two cyclic groups $H_{1}=\langle(1,2,3,4)\rangle$ and $H_{2}=\langle(5,6,7)\rangle$
In this section we denote by $z(i)$ the coordinates of $z$ that are active in the i-th cyclic group in the direct product $H_{1}\times\ldots{}\times H_{d}$.
The following theorem states that if a permutation group $G\leqslant S_{n}$ is a direct product of other permutation groups $G^{i}$, (that is, permutation groups $G^{i}$ acting on disjoint subsets of coordinates in the usual way), then the core set of $G$ is also a Cartesian product.
Theorem 5.1
(5, , Theorem 8) Let $G=\prod_{i=1}^{d}G^{i}$, $G^{i}\leqslant S_{n}$. Then
$\mathrm{core}(G)=\prod_{i=1}^{d}\mathrm{core}(G^{i}).$
Proof
The product structure of $G$ induces a decomposition of $\mathbb{R}^{n}$ into a Cartesian product of pairwise orthogonal coordinate subspaces $\oplus_{i=1}^{d}X_{i}=\mathbb{R}^{n}$. Thus, we can write every point $z\in\mathbb{R}^{n}$ as $z=\oplus_{i=1}^{d}z(i)$. The claim of the theorem follows immediately from $\mathrm{conv}(G_{z})=\prod_{i=1}^{d}\mathrm{conv}(G^{i}_{z(i)})$.
We are concerned with various subgroups of the direct product $\prod_{j=1}^{d}H_{j}$, where the cyclic groups $H_{j}=\langle h_{j}\rangle$ are generated by disjoint cycles $h_{1},\ldots,h_{d}$ in $S_{n}$. Let $k_{j}$ denote the length (period) of the cycle $h_{j}$. Thus, if $g=h_{1}\cdots h_{d}$ is the complete factorization of the permutation $g\in S_{n}$ into disjoint (possibly trivial) cycles, then $k_{1}+\cdots k_{d}=n$.
Note that we can use the same $\widehat{T}(c)$ as in Lemma 1 for each cycle of the direct product groups. In other words, let $c=\bigoplus_{i=1}^{d}c(i)\in\mathbb{Z}^{n}$. When required, we will suppose that the vector $c$ has the property that the matrices $\operatorname{Cir}(c(j))$ are invertible, $1\leqslant j\leqslant d$. Lemma 1 then tells us the corresponding inverse matrices are $\operatorname{Cir}(\widehat{T}(c(j)))$. Then a point $x=\bigoplus_{i=1}^{d}x(i)\in\mathbb{R}^{n}$ is in $\mathrm{conv}(G_{c})$ if and only if
$x(1)=\operatorname{Cir}(c(1))\lambda^{1}$ $\Rightarrow\lambda^{1}=\operatorname{Cir}(\widehat{T}(c(1)))x(1)\,,$
$\vdots$
$x(d)=\operatorname{Cir}(c(d))\lambda^{d}$ $\Rightarrow\lambda^{d}=\operatorname{Cir}(\widehat{T}(c(d)))x(d)\,.$
Furthermore, by Theorem 5.1 we can apply our new constraints on each subspace.
We denote by $l^{j}_{i}$ the sub-layer of active variables of an integer point in the $j$-th cyclic group and in the $i$-th equivalence class, and denote by $\widehat{E}^{j}_{i}$ the corresponding projected essential set of each sub-layer $l_{i}^{j}.$
Similarly to the previous section, all equivalence classes of all sub-layers must be checked. By Lemma 3, Remark 5 and Lemma 7, in each cyclic group $H_{i}$ we need to search in $k_{i}$ sub-layers. Then there are $k_{1}\times\ldots{}\times k_{d}$ possibilities for sub-layers of an integer point. For example suppose $G=\langle(1,2,3)\rangle\times\langle(4,5)\rangle\leqslant S_{5}$. Then we have decomposition $\mathbb{Z}^{n}=X_{1}\times X_{2}$ where $z(1)=(z_{0},z_{1},z_{2})$ and $z(2)=(z_{3},z_{4})$. So we have the following possibilities for sub-layers of a feasible solution $z$
$l^{1}_{1}$ $=3q_{1}+1,\;\;l^{2}_{1}=2q_{2}+1,$
$l^{1}_{2}$ $=3q_{1}+2,\;\;l^{2}_{1}=2q_{2}+1,$
$l^{1}_{3}$ $=3q_{1}+3,\;\;l^{2}_{1}=2q_{2}+1,$
$l^{1}_{1}$ $=3q_{1}+1,\;\;l^{2}_{2}=2q_{2}+2,$
$l^{1}_{2}$ $=3q_{1}+2,\;\;l^{2}_{2}=2q_{2}+2,$
$l^{1}_{3}$ $=3q_{1}+3,\;\;l^{2}_{2}=2q_{2}+2,$
where $q_{1}$ and $q_{2}$ are arbitrary integers.
Consider a cycle $h_{j}$. Since its length is $k_{j}$, a complete set of residues modulo $k_{j}$ is the set $[k_{j}]:=\{1,\ldots,k_{j}\}$. We require their Cartesian product
[…]
For each element $(t_{1},\ldots{},t_{d})$ of $K\subset\mathbb{Z}^{n}$, a subproblem of type $Q_{1}$ is solved, where $t_{j},\;j=1,\ldots{},d$, is the sub-layer of active variables of cycle $h_{j}.$
Remark 8
In the direct product group $G=\prod_{i=1}^{d}H_{i}$, each sub-layer has its own projected essential set. In step 4 of Algorithm 2 the feasibility of pre-images of integer points of the projected essential sets is checked. In the direct product case we can check the feasibility of these pre-images corresponding to each cycle separately. If $s_{j}:=\sum_{i=1}^{k_{j}}|\widehat{E}^{j}_{i}|,$ then $\sum_{j=1}^{d}s_{j}$ subproblems must be solved.
On the other hand if adding constraints for one cycle is not enough, we can check for simultaneous feasibility. Let $(t_{1},\dots,t_{d})\in K$ be a vector of residues (where $K$ is defined above). Given points $z(j)\in\widehat{E}^{j}_{t_{j}}$, $j=1\ldots d$, we can check simultaneous feasibility of $z(1),\ldots,z(d)$. In this case, $\Pi_{j=1}^{d}s_{j}$ subproblems must be solved.
For example if $G=C_{5}\times C_{8}$ and a universal core point in the projected essential set $\widehat{E}^{1}_{2}$ is $(1,1,0,0,0)$, by Remark 7, we add the following constraints to the problem:
$x_{1}$ $=x_{0}\,,$
$x_{2}$ $=x_{0}-1\,,$
$x_{3}$ $=x_{0}-1\,,$
$x_{4}$ $=x_{0}-1\,.$
We can also check all possible combinations of integer points in the projected essential sets. For example if $G=\langle(1,2,3,4)\rangle\times\langle(5,6,7)\rangle\leqslant S_{7}$ , for $l_{2}^{1}$ and $l_{1}^{2}$ let $\widehat{E}_{2}^{1}=\{(1,1,0,0),(1,0,1,0)\}$ and $\widehat{E}_{1}^{2}=\{(1,0,0),(2,-1,0)\}$. Then all possible combinations of $\widehat{E}_{2}^{1}$ and $\widehat{E}_{1}^{2}$ are as below:
$(1,1,0,0,1,0,0),(1,1,0,0,2,-1,0),(1,0,1,0,1,0,0),(1,0,1,0,2,-1,0).$
Now we can generalize Algorithm 2 for direct product groups as shown in Algorithm 3. For simplicity we present the version which tests the essential points individually for feasibility; the modification of the loop at step 3 for simultaneous testing is straightforward.
Algorithm 3 Maximization for ILPs with direct product cyclic symmetry
maximization ILP $P$with symmetry group $G=\Pi_{j=1}^{d}H_{j}$, bounded relaxation, objective function $f$. The optimal objective value, or $\infty$- ∞if $P$is infeasible.
1:$f^{*}\leftarrow-\infty$
[…]
4:Construct and solve subproblem $P^{z}_{t}$ of type $Q_{3}$ by adding constraints described in Remark 7 to $P$ for point $z$.
5:$f^{*}\leftarrow\max\{f^{*},\operatorname{OPT}_{\infty}(P^{z}_{t})\}$EndFor
[…]
9:$f^{*}\leftarrow\max\{f^{*},\operatorname{OPT}_{\infty}(P_{t})\}$EndFor
[…]
Remark 9
Compared to Algorithm 2, both the number of subproblems and their difficulty are potential bottlenecks in Algorithm 3. While the steps are similar to the Algorithm 2, the running time of each subproblem is longer. In particular in each subproblem of step 4 we either have constraints acting on a small set of coordinates (if we check each subgroup individually), or a very large number of subproblems (if we check for simultaneous feasibility).
6 New Constraints For Permutation Groups
In this section we use Algorithms 3, 2 or 1 for a symmetric ILP with some permutation group as its symmetry group. First we define subdirect product groups and then classify the permutation groups with respect to their generators.
Definition 15
Let $H_{1}$ and $H_{2}$ be groups. A subdirect product of $H_{1}$ and $H_{2}$ is a subgroup $H$ of the external direct product $H_{1}\times H_{2}$ such that the projection from $H$ to either direct factor is surjective. In other words, if $p_{1}:H_{1}\times H_{2}\to H_{1}$ is given by $(h_{1},h_{2})\mapsto h_{1}$ and $p_{2}:H_{1}\times H_{2}\to H_{2}$ is given by $(h_{1},h_{2})\mapsto h_{2}$, then $p_{1}(H)=H_{1}$ and $p_{2}(H)=H_{2}$. Subdirect products of more than two groups are obtained naturally by iterating this construction.
Example 3
The group $\langle(1,2)(3,4,5,6)\rangle$ is a subdirect product of $\langle(1,2)\rangle$ and $\langle(3,4,5,6)\rangle$ and is a subgroup of the direct product $\langle(1,2)\rangle\times\langle(3,4,5,6)\rangle$.
Remark 10
Suppose $g\in G\leqslant S_{n}$ is factored as a product of disjoint cycles. These cycles themselves may or may not be elements of $G$. But Lemma 5 shows that for exploring feasible integer points of a symmetric polyhedron considering some of the cycles is enough. In other words, we can make our constraints for a few cycles and keep other variables non-active.
Note that our constraints are still valid for a subdirect product of disjoint cycles, since in each sub-problem we remove universal or atom points (by generating constraints from their projections) from the orbit polytope of active variables of each cycle. So Algorithms 2 and 3 work for this case as well.
Now we classify permutation groups with respect to their generators.
Disjoint Cycles
If $G$ is a group generated by disjoint cycles $h_{1},\ldots,h_{d}$, then $G$ is the direct product of $H_{1},..,H_{d}$ (Lemma 7). In this case Algorithm 3 can be applied.
Example 4
The group $G=\langle(1,2,3,4),(6,8,9),(5,7,12,11)\rangle$ is the direct product of the three cyclic groups $H_{1}=\langle(1,2,3,4)\rangle$, $H_{2}=\langle(6,8,9)\rangle$, $H_{3}=\langle(5,7,12,11)\rangle.$
Product of disjoint cycles
If the given set of generators of a group $G$ is a single permutation which is the product of two or more disjoint cycles, then $G$ is a subdirect product of the corresponding cyclic groups. In this case by Remark 10, Algorithm 3 can be applied.
Example 5
The subdirect product group $G=\langle(1,2,5,3)(6,10,11)(9,7,12,8)\rangle$ is a subgroup of the direct product of three cyclic groups $H_{1}=\langle(1,2,5,3)\rangle$, $H_{2}=\langle(6,10,11)\rangle$, $H_{3}=\langle(9,7,12,8)\rangle.$
Combination of two of the above cases
If the generators of $G$ are products of various numbers of disjoint cycles, then Algorithm 3 can be applied.
Example 6
Let $G=\langle(1,2,3,4)(5,6,12,13),(7,8,9,10)\rangle$, then $G$ is a subgroup of the direct product of three cycles $H_{1}=\langle(1,2,3,4)\rangle$, $H_{2}=\langle(5,6,12,13)\rangle$, $H_{3}=\langle(7,8,9,10)\rangle$.
Non-disjoint cycles
If the given set of generators of $G$ are not disjoint cycles, we find a subgroup of $G$ where all generators are disjoint. Then it falls into one of the previous three cases. Note that for finding the biggest cyclic subgroup we can find representatives of the conjugacy classes (e.g. by using GAP GAP4 ) and choose the biggest cycle.
Example 7
Let $G=\langle(1,2,3,4,5),(7,5,8,10,11)(12,13,2,14)\rangle$ then $H_{1}=\langle(1,14,12,13,2,3,4,8,10,11,7,5)\rangle$ is a cyclic subgroup and since 9 and 6 are fixed, Algorithm 2 can be applied.
7 Computational Experiments
In order to test the efficiency of our algorithms, in this section we create some symmetric integer linear programs that are hard to solve with standard solvers. For this purpose, we can use the orbit polytopes of core points. As the orbit polytope of a core point contains no integer points aside from the vertices, if we can cut off these vertices then the integer program corresponding to this polytope will be infeasible.
Infeasible problems are typically hard for branch-and-bound algorithms because there is no chance of early success. The goal of the integer feasibility problems is to find an integer point in a polyhedron $P$ or decide that no such point exists. Aaarden and Lenstra presented some difficult integer feasibility problems AL04 ; GZ02 whose relaxations are simplices. Our techniques are not suitable for the examples of (AL04, , Table 1) because these simplices lack symmetries in the sense of Remark 1. Symmetric instances are interesting in their own right because equivalent partial solutions can blow up the branch-and-bound-tree (cf. PR2019 ). In order to construct symmetric examples that are still integer infeasible, we start with lattice-free symmetric simplices and cut off the (integer) vertices; a similar technique is used in 2 .
The convex hull of the orbit of a core point under a cyclic group with an invertible $n\times n$ circulant matrix is a simplex with dimension $n-1$. We say an integer point has a globally minimal projection (with respect to some invariant subspace $V$) if
$\|z|_{V}\|\leqslant\|z^{\prime}|_{V}\|\text{ for all }z^{\prime}\in\mathrm{aff}(G_{z})\bigcap\mathbb{Z}^{n}.$
For a primitive group the following theorem shows that the corresponding orbit polytope is often a simplex.
Theorem 7.1
(2, , Theorem 5.37) Let $G\leqslant S_{n}$ be primitive and let $V\leqslant\mathbb{R}^{n}$ be a rational invariant subspace. If $e_{0}=(1,0,0,\ldots{},0)$ has globally minimal projection onto $V$, then there are infinitely many core points in layer one. The corresponding orbit polytopes are simplices.
Based on Theorem 7.1 the core point used to define the instances is calculated as $p+\sigma w$ for some universal core point $p$ and direction $w$ orthogonal to the fixed space. We refer to the scalar $\sigma$ as the scale factor or just scale. The groups used for these experiments are the primitive groups with GAP identifiers $(15,2)$, $(21,2)$, $(45,1)$ (all three of them have rational invariant subspaces). Table 1 gives more details about symmetry groups of these instances.
Preliminary experiments indicated that for these problems, there was little benefit to using Algorithm 3 instead of Algorithm 2. In Section 3.3 we presented an example where the subproblems were solved using Knitro knitro . While Knitro performs well in our setting, we wanted to ensure our results were replicable with open source software. For this reason we have implemented Algorithm 2 using PySCIPOpt PySCIPOpt . In order to simplify the nonlinear subproblems $Q_{1}$, we have only implemented the case of single element essential sets, using (13).
We use Algorithm 2, CPLEX, Gurobi and SCIP SCIP to solve P2, P3, and P4 and compare the results. We also checked that as expected the analogous feasible instances (simplices) are generally easy for all solvers, and that Algorithm 2 finds the expected solution. All experiments in this section were carried out a Dell PowerEdge R640 server with a Xeon Gold 6248 (2.5GHz) processor and 384GB of RAM. The operating system was Debian Linux 12.10 (amd64). CPLEX was version 22.1.1 and Gurobi was version 12.0.1. We used SCIP 9.2.1 and PySCIPOpt 5.3.0. All results are the mean of 3 runs. We use a 1 hour time limit for all runs.
Table 1: ILP-feasibility problems.
Name Symmetry group Generators of the group
P2 Primitive group (15,2) $(1,15,7,5,12)(2,9,13,14,8)(3,6,10,11,4)$, $(1,4,5)(2,8,10)(3,12,15)(6,13,11)(7,9,14)$
P3 Primitive group (21,2) (1,7,12,16,19,21,6)(2,8,13,17,20,5,11)(1,7,12,16,19,21,6)(2,8,13,17,20,5,11$\cdot$( 1 , 7 , 12 , 16 , 19 , 21 , 6 ) ( 2 , 8 , 13 , 17 , 20 , 5 , 11 ) ⋅ $(3,9,14,18,4,10,15)$, $(4,6,5)(9,11,10)(13,15,14)(16,18,17)(19,20,21)$
P4 Primitive group (45,1) (1,2,7)(3,11,27)(4,14,31)(5,18,32)(6,20,36)(1,2,7)(3,11,27)(4,14,31)(5,18,32)(6,20,36$\cdot$( 1 , 2 , 7 ) ( 3 , 11 , 27 ) ( 4 , 14 , 31 ) ( 5 , 18 , 32 ) ( 6 , 20 , 36 ) ⋅ (8,24,39)(9,25,28)(10,26,42)(12,15,16)(13,30,40)(8,24,39)(9,25,28)(10,26,42)(12,15,16)(13,30,40$\cdot$( 8 , 24 , 39 ) ( 9 , 25 , 28 ) ( 10 , 26 , 42 ) ( 12 , 15 , 16 ) ( 13 , 30 , 40 ) ⋅ (17,19,21)(22,35,44)(23,33,29)(34,43,37)(17,19,21)(22,35,44)(23,33,29)(34,43,37$\cdot$( 17 , 19 , 21 ) ( 22 , 35 , 44 ) ( 23 , 33 , 29 ) ( 34 , 43 , 37 ) ⋅ $(38,45,41)$, (1,3,5,6,7,22,13,23)(2,8,9,10)(1,3,5,6,7,22,13,23)(2,8,9,10$\cdot$( 1 , 3 , 5 , 6 , 7 , 22 , 13 , 23 ) ( 2 , 8 , 9 , 10 ) ⋅ (4,15,16,17,14,21,19,12)(4,15,16,17,14,21,19,12$\cdot$( 4 , 15 , 16 , 17 , 14 , 21 , 19 , 12 ) ⋅ (11,28,29,38,44,25,20,37)(11,28,29,38,44,25,20,37$\cdot$( 11 , 28 , 29 , 38 , 44 , 25 , 20 , 37 ) ⋅ (18,33,34,24,40,36,41,39)(18,33,34,24,40,36,41,39$\cdot$( 18 , 33 , 34 , 24 , 40 , 36 , 41 , 39 ) ⋅ $(26,43,35,32,42,45,27,30)$, (1,4)(3,12)(5,19)(6,21)(7,14)(8,10)(11,20)(1,4)(3,12)(5,19)(6,21)(7,14)(8,10)(11,20$\cdot$( 1 , 4 ) ( 3 , 12 ) ( 5 , 19 ) ( 6 , 21 ) ( 7 , 14 ) ( 8 , 10 ) ( 11 , 20 ) ⋅ (13,16)(15,23)(17,22)(18,33)(24,41)(25,28)(13,16)(15,23)(17,22)(18,33)(24,41)(25,28$\cdot$( 13 , 16 ) ( 15 , 23 ) ( 17 , 22 ) ( 18 , 33 ) ( 24 , 41 ) ( 25 , 28 ) ⋅ $(26,43)(27,32)(29,44)(30,35)(34,39)(36,40)(42,45)$
Figure 2: Comparison of elapsed time for problem P2. Asterisks indicate timeout.
1111222233334444555566667777888899991010101011111111121212121313131314141414151515151616161617171717181818181919191920202020005005005005001,00010001{,}0001 , $0002{,}$ $0003{,}$ 0003,60036003{,}6003 , 600ScaleElapsed Time (s)Algorithm 2CPLEXGurobiSCIP
Table 2: Solution times (seconds) and memory usage (kilobytes) for P2.
Solver Scale Elapsed Time (s) Kernel (s) User (s) Memory (kB)
alg2 11 46.66 0.26 45.56 93492
cplex 11 468.63 0.01 468.56 25168
gurobi 11 35.55 0.01 35.50 15728
scip 11 47.36 0.12 47.23 44688
alg2 12 103.73 0.29 102.45 110489
cplex 12 126.68 0.00 126.62 21583
gurobi 12 0.52 0.00 0.48 15083
scip 12 73.86 0.23 73.40 85620
alg2 13 48.11 0.28 46.93 99559
cplex 13 1166.74 0.03 1166.61 47459
gurobi 13 310.69 0.00 310.64 17308
scip 13 16.13 0.14 16.00 49881
alg2 14 127.67 0.37 126.24 117356
cplex 14 1146.67 0.03 1146.55 42931
gurobi 14 360.81 0.01 360.75 19704
scip 14 50.03 0.22 49.68 74865
alg2 15 66.69 0.31 65.50 104793
cplex 15 1097.42 0.01 1097.34 21613
gurobi 15 298.17 0.01 298.12 17077
scip 15 129.70 0.40 128.77 155315
alg2 16 689.64 1.03 686.77 292491
cplex 16 1886.85 0.05 1925.08 54880
gurobi 16 646.88 0.02 646.81 21843
scip 16 163.80 0.12 2.76 49144
alg2 17 1885.64 2.52 1879.69 743260
cplex 17 2610.94 0.13 2610.65 104028
gurobi 17 776.52 0.02 776.43 19444
scip 17 160.27 0.51 159.07 128180
alg2 18 1275.66 1.34 1271.31 458885
gurobi 18 775.60 0.01 775.51 20956
scip 18 187.94 0.54 186.53 111533
alg2 19 725.15 1.18 722.11 446423
gurobi 19 1631.70 0.01 1631.59 22445
scip 19 264.65 0.64 262.91 179401
alg2 20 3475.37 4.22 3465.94 1312812
gurobi 20 2205.33 0.05 2205.17 37880
scip 20 233.04 0.77 231.14 197280
We first discuss problem P2. As shown in Figure 2, for scales 1 to 10, all solvers complete in less than 5 minutes. Table 2 shows a comparison of time and memory use for scales 11 through 20. The built-in symmetry handling of SCIP (not used by Algorithm 2) performs well on these problems. For scales 11-20 Algorithm 2 is generally faster than CPLEX and Gurobi, and in particular completes all 20 scales within the 1 hour time limit, while CPLEX times out for scales 18, 19, and 20.
Table 3: Solution times and memory usage for P3.
Solver Scale Elapsed Time (s) Kernel (s) User (s) Memory (kB)
alg2 1 9.24 0.17 9.17 74957
cplex 1 0.78 0.003 0.77 16981
scip 1 2.00 0.11 1.99 28257
gurobi 1 5.96 0.007 5.95 15997
alg2 2 149.42 0.29 149.22 117337
cplex 2 20.95 0.000 20.95 17921
scip 2 5.63 0.11 5.63 29745
gurobi 2 91.47 0.003 91.46 16329
alg2 3 363.08 0.377 362.77 164063
cplex 3 108.43 0.010 108.41 21687
scip 3 191.34 0.31 191.13 78953
gurobi 3 1075.75 0.017 1075.72 23979
alg2 4 1074.83 0.63 1074.26 219821
cplex 4 TO TO TO -
scip 4 TO TO TO -
gurobi 4 2342.67 0.033 2342.59 43643
alg2 5 2036.46 1.357 2035.11 369835
cplex 5 TO TO TO -
scip 5 TO TO TO -
gurobi 5 4350.08 0.06 4349.00 75039
Table 3 presents the performance of Algorithm 2, CPLEX, SCIP, and Gurobi on problem P3 for scales 1 to 5. The solvers did not solve the problem for scales 6 to 10 in less than 1 hour. Algorithm 2 and Gurobi are the only solvers that complete for 5 scales — two more than CPLEX and SCIP, which solve instances only up to scale 3 before exceeding the 1 hour time limit.
Table 4: Solution times and memory usage for P4.
Solver Scale Elapsed Time (s) Kernel (s) User (s) Memory (kB)
alg2 1 81.54 0.31 81.23 94292
cplex 1 2.97 0.003 2.94 18261
scip 1 7.25 0.11 7.21 41339
gurobi 1 6.64 0.00 6.61 16517
alg2 2 1081.36 0.64 1080.66 209033
cplex 2 103.02 0.007 103.01 24896
scip 2 65.00 0.29 64.81 173423
gurobi 2 751.74 0.02 751.71 19220
alg2 3 TO TO TO -
cplex 3 1813.18 0.04 1813.06 44367
scip 3 TO TO TO -
gurobi 3 TO TO TO -
Table 4 compares the performance of Algorithm 2, CPLEX, SCIP, and Gurobi on the infeasible ILP problem P4 for scales 1 to 3. All solvers complete for scales 1 and 2. At scale 3, only CPLEX completes, while the others time out. For scales 4 to 10, all solvers time out. Among the four, Gurobi demonstrates the best performance at smaller scales in terms of both runtime and memory usage. SCIP also shows strong performance on scale 2 but with higher memory usage. These instances seem challenging for all tested solvers.
8 Conclusions
In this paper we introduced some new techniques for solving symmetric linear programs based on deriving nonlinear constraints from the symmetry group of the formulation. Since these constraints depend only on the symmetry group, the same constraints can be re-used for many problems. The practical benefits of using a nonlinear solver in order have a smaller search space need more evaluation, but at least on the artificial instances in Section 7 they show some promise, allowing the solution of instances not solvable within 1 hour on the same hardware using commercial MILP solvers. We think our methods may be useful for problems with a large enough cyclic subgroup in their symmetry group, particularly those where determining integer feasibility is a challenge. From a theoretical point of view, these techniques show that core point techniques are not limited to groups where the number of (non-equivalent) core-points is finite.
Acknowledgements
The authors wish to thank two anonymous referees for many helpful suggestions, including the simplified proof of Lemma 2 that is given here. They would also like to thank Barry Monson, Branimir Ćaćic and Nicholas Touikan for helpful feedback on earlier versions of the paper.